# Integrable open elliptic Toda chain with boundaries

### v1

@@ 2 Lax matrices in factorized form
This is an intertwining matrix entering the IRF-Vertex correspondence [3] in quantum statistical models. We use the standard notations for Jacobi theta functions. From here on it is assumed that all elliptic functions depend on $\tau$ unless otherwise stated explicitly (like in (2)). Let us introduce a notation for a set of $2\times 2$ matrices:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\qquad P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})}\end{array}$ (3)
Direct calculation yields expression for $L^{[k,m]}(z)$ in terms of the elliptic Kronecker function $\phi(z,x)$:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,\quad\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (4)
where
$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\ \ b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (5)
@@
The Lax matrices for the elliptic Toda chain can be written in several different ways. First possibility is to define the monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ as
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,,}\end{array}$ (6)
where we use the Lax matrices (4). Another possibility is to use the modified Lax representation [16, 10]:
$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)\,,\qquad{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (7)
Then ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Explicitly,
$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (8)
Using the identity $\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)$ we have
$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)\,.}\end{array}$ (9)
Then it is easy to see that $\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)$ provides the Hamiltonian (1).
@@ 3 Gauge equivalence with XYZ chain
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}}\end{array}$ (10)
Therefore, the gauge transformed monodromy matrix ${\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})$ takes the form
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)\,,\qquad{\mathcal{L}}^{k}(z)=g(z,{\bf{q}}_{k})e^{P^{k}/2c}g^{-1}(z,{\bf{q}}_{k})}\end{array}$ (11)
@@
It was explained in [10] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [12, 4]. Namely,
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,\qquad\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,,}\end{array}$ (12)
where $\sigma_{a}$ are the Pauli matrices ($\sigma_{0}$ is $2\times 2$ identity matrix). The variables ${\bf S}^{a}_{k}$ are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as
$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)}\end{array}$ (13)
for all a=1,,na=1,...,n and $k=1,2,3$. The Poisson brackets between ${\bf S}_{k}^{a}$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [11]
@@
$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (14)
@@
where the half-periods $\omega_{j}$
@@
are numerated according to numeration of the Pauli matrices. The Sklyanin algebra (14) is generated by the classical quadratic exchange relation
$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z),{\mathcal{L}}_{2}^{b}(w)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z){\mathcal{L}}_{2}^{a}(w),r_{12}(z-w)]\,,\qquad r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (16)
The representation (13) for the variables ${\bf S}^{a}_{k}$ corresponds to special values of the Casimir functions ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$:

### v3

# arXiv:2602.13903v3[nlin.SI] 28 Feb 2026

@@ 2 Lax matrices in factorized form
This is an intertwining matrix entering the IRF-Vertex correspondence [4] in quantum statistical models. Its geometric interpretation and applications to interrelations between classical integrable systems can be found in [9, 18, 3]. We use the standard notations for Jacobi theta functions. From here on it is assumed that all elliptic functions depend on $\tau$ unless otherwise stated explicitly (like in (2)). Let us introduce a notation for a set of $2\times 2$ matrices:
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=-g^{-1}(z,{\bf{q}}_{k})g(z,{\bf{q}}_{m})e^{P^{m}/2c}\,,\quad k,m=1,...,n\,,}\end{array}$ (3)
where $P^{m}$ are diagonal $2\times 2$ matrices of particles momenta:
$\begin{array}[]{c}\displaystyle{P^{m}={\rm diag}({\bf{p}}_{m},-{\bf{p}}_{m})\,,\quad m=1,...,n\,.}\end{array}$ (4)
Direct calculation yields expression for $L^{[k,m]}(z)$
$\begin{array}[]{c}\displaystyle{L^{[k,m]}(z)=\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\\ \\ {\phi(z,-{\bf{q}}_{k}-{\bf{q}}_{m})b^{[k,m]}_{1}}&{\phi(z,-{\bf{q}}_{k}+{\bf{q}}_{m})b^{[k,m]}_{2}}\end{array}\right)\,,}\end{array}$ (5)
where $\phi(z,x)$ is the elliptic Kronecker function
$\begin{array}[]{c}\displaystyle{\phi(z,u)=\frac{\theta_{1}^{\prime}(0)\theta_{1}(z+u)}{\theta_{1}(z)\theta_{1}(u)}}\end{array}$ (6)
and $b^{[k,m]}_{1}$, $b^{[k,m]}_{2}$ are the following functions:
$\begin{array}[]{c}\displaystyle{b^{[k,m]}_{1}=\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{{{\bf{p}}_{m}}/{2c}}\,,}\\ \\ \displaystyle{b^{[k,m]}_{2}=-\frac{\theta_{1}({\bf{q}}_{m}-{\bf{q}}_{k})\theta_{1}({\bf{q}}_{m}+{\bf{q}}_{k})}{\theta_{1}^{\prime}(0)\theta_{1}(2{\bf{q}}_{m})}\,e^{-{{\bf{p}}_{m}}/{2c}}\,.}\end{array}$ (7)
@@
The Lax matrices for the elliptic Toda chain can be written in several different ways. First possibility is to use $L^{[a-1,a]}(z)$, a=1,,na=1,...,n from the definition (5). The monodromy matrix $T^{\hbox{\tiny{eToda}}}(z)$ takes the form [13]:
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=L^{[n,1]}(z)L^{[1,2]}(z)L^{[2,3]}(z)\dots L^{[n-1,n]}(z)\,.}\end{array}$ (8)
Here $z$ is the spectral parameter. Being written in this from, the elliptic Toda chain is a particular case of ${\rm GL}_{N}$ elliptic Ruijsenaars chain on $n$ sites introduced in [20]. Its classical $r$-matrix structure was described in [12]. More precisely, ${\rm GL}_{2}$ elliptic Ruijsenaars chain depends on the additional deformation parameter $\eta$, and the elliptic Toda chain arises when $\eta=0$, while for $\eta\neq 0$ it reproduces another known integrable model – the elliptic Ruijsenaars-Toda chain [^1] In fact, the elliptic Ruijsenaars-Toda chain contains $n$ parameters $\eta_{a}$, a=1,,na=1,...,n. That is, the elliptic Ruijsenaars-Toda chain with $\eta_{a}=0$ for all $a$ yields the elliptic Toda chain, and the elliptic Ruijsenaars-Toda chain with pairwise equal $\eta_{a}=\eta_{b}$ corresponds to ${\rm GL}_{2}$ elliptic Ruijsenaars chain (with reduction to the center of mass frames at each site). Details of interrelations between models can be found in [13]. described in [2].
Alternatively, one can use the modified Lax representation:
$\begin{array}[]{c}\displaystyle{{\bf T}^{\hbox{\tiny{eToda}}}(z)={\bf L}^{1}(z){\bf L}^{2}(z)\dots\bf L^{n}(z)}\end{array}$ (9)
with
$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{b^{[a-1,a]}_{1}+b^{[a-1,a]}_{2}}\,L^{[a-1,a]}(z)\,.}\end{array}$ (10)
Explicitly, we have
$\begin{array}[]{c}\displaystyle{{\bf L}^{a}(z)=\frac{1}{2}\left(\begin{array}[]{cc}{\phi(z,{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\\ \\ {\phi(z,-{\bf{q}}_{a-1}-{\bf{q}}_{a})\displaystyle{\frac{e^{{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}&{-\phi(z,-{\bf{q}}_{a-1}+{\bf{q}}_{a})\displaystyle{\frac{e^{-{\bf{p}}_{a}/2c}}{\sinh({\bf{p}}_{a}/2c)}}}\end{array}\right)\,.}\end{array}$ (11)
The classical transfer-matrices in these two descriptions are related as ${\bf T}^{\hbox{\tiny{eToda}}}(z)={T}^{\hbox{\tiny{eToda}}}(z)e^{-H^{\hbox{\tiny{eToda}}}/2}$. Using the identity
$\begin{array}[]{c}\displaystyle{\phi(z,u)\phi(z,-u)=\wp(z)-\wp(u)}\end{array}$ (12)
for the function (6) we have
$\begin{array}[]{c}\displaystyle{\det{\bf L}^{a}(z)=\frac{1}{4}\frac{1}{\sinh^{2}({\bf{p}}_{a}/2c)}\Big(\wp({\bf{q}}_{a-1}-{\bf{q}}_{a})-\wp({\bf{q}}_{a-1}+{\bf{q}}_{a})\Big)}\end{array}$ (13)
for ${\bf L}^{a}(z)$ (11). Notice that it is independent of spectral parameter $z$. Then it is easy to see that the expression
$\begin{array}[]{c}\displaystyle{\log\det{\bf T}^{\hbox{\tiny{eToda}}}(z)=\sum\limits_{a=1}^{n}\log\det{\bf L}^{a}(z)}\end{array}$ (14)
provides the original Hamiltonian (1).
Details of the above results can be found in [20, 13]. Let us also mention that we use the factorization formulae of type (3). Originally, this type representation of the Lax matrices was observed by K. Hasegawa in [6] at quantum level for the Ruijsenaars-Schneider model. See also a review [18] and references therein.
@@ 3 Gauge equivalence with XYZ chain
$\begin{array}[]{c}\displaystyle{T^{\hbox{\tiny{eToda}}}(z)=}\\ \\ \displaystyle{=(-1)^{n}g^{-1}(z,{\bf{q}}_{n})g(z,{\bf{q}}_{1})e^{P^{1}/2c}g^{-1}(z,{\bf{q}}_{1})g(z,{\bf{q}}_{2})e^{P^{2}/2c}...g^{-1}(z,{\bf{q}}_{n-1})g(z,{\bf{q}}_{n})e^{P^{n}/2c}\,.}\end{array}$ (15)
Therefore, the gauge transformed monodromy matrix
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=g(z,{\bf{q}}_{n})T^{\hbox{\tiny{eToda}}}(z)g^{-1}(z,{\bf{q}}_{n})}\end{array}$ (16)
takes the form
$\begin{array}[]{c}\displaystyle{{\mathcal{T}}(z)=(-1)^{n}{\mathcal{L}}^{1}(z){\mathcal{L}}^{2}(z)\dots\mathcal{L}^{n}(z)}\end{array}$ (17)
with the Lax matrices
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=g(z,{\bf{q}}_{a})e^{P^{a}/2c}g^{-1}(z,{\bf{q}}_{a})\,,\quad a=1,...,n\,.}\end{array}$ (18)
@@
It was explained in [13] that ${\mathcal{L}}^{a}(z)$ are the Lax matrices of the XYZ chain, and ${\mathcal{T}}(z)$ is its standard monodromy matrix [15, 5]. Namely, the matrices ${\mathcal{L}}^{a}(z)$ are represented in the form:
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}\,,}\end{array}$ (19)
where $\sigma_{a}$ are the Pauli matrices
@@
$\begin{array}[]{c}\displaystyle{\sigma_{0}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{1}\end{array}\right)\,,\quad\sigma_{1}=\left(\begin{array}[]{cc}{0}&{1}\\ {1}&{0}\end{array}\right)\,,\quad\sigma_{2}=\left(\begin{array}[]{cc}{0}&{-\imath}\\ {\imath}&{0}\end{array}\right)\,,\quad\sigma_{3}=\left(\begin{array}[]{cc}{1}&{0}\\ {0}&{-1}\end{array}\right)}\end{array}$ (20)
and
$\begin{array}[]{c}\displaystyle{\varphi_{1}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{4}(z)}{\theta_{1}(z)\theta_{4}(0)}\,,\qquad\varphi_{2}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{3}(z)}{\theta_{1}(z)\theta_{3}(0)}\,,\qquad\varphi_{3}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{2}(z)}{\theta_{1}(z)\theta_{2}(0)}\,,}\end{array}$ (21)
or, in the short form
$\begin{array}[]{c}\displaystyle{\varphi_{k}(z)=\frac{\theta_{1}^{\prime}(0)\theta_{5-k}(z)}{\theta_{1}(z)\theta_{5-k}(0)}\,,\quad k=1,2,3\,.}\end{array}$ (22)
The variables ${\bf S}^{a}_{k}$ entering (19) are expressed through ${\bf{p}}_{a}$ and ${\bf{q}}_{a}$ as follows:
$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}=\frac{1}{2}\Big(e^{{\bf{p}}_{a}/2c}+e^{-{\bf{p}}_{a}/2c}\Big)\,,}\\ \\ \displaystyle{{\bf S}_{k}^{a}=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a})}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad k=1,2,3}\end{array}$ (23)
for all a=1,,na=1,...,n. This is a particular case of a more general formulae [^2] In the continuous ($n\to\infty$) and non-relativistic ($c\to\infty$) limit the XYZ chain turns into 1+1 XYZ Landau-Lifshitz model. It is gauge equivalent to 1+1 field analogue of 2-body Calogero-Moser model [9, 3]. In this respect, the elliptic Toda chain in the continuous non-relativistic limit provides this 1+1 field analogue of 2-body Calogero-Moser model with zero coupling constant.:
$\begin{array}[]{c}\displaystyle{{\mathcal{L}}^{a}(z,\eta)=\sigma_{0}{\bf S}^{a}_{0}+\sum\limits_{k=1}^{3}\sigma_{k}\varphi_{k}(z){\bf S}^{a}_{k}(\eta)\,,}\end{array}$ (24)
$\begin{array}[]{c}\displaystyle{{\bf S}_{0}^{a}(\eta)=\frac{1}{2}\Big(\frac{\theta_{1}(2{\bf{q}}_{a}-\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}+\frac{\theta_{1}(2{\bf{q}}_{a}+\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,}\\ \\ \displaystyle{{\bf S}_{k}^{a}(\eta)=\frac{1}{2}\frac{\theta_{5-k}(0)}{\theta_{1}^{\prime}(0)}\Big(\frac{\theta_{5-k}(2{\bf{q}}_{a}-\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{{\bf{p}}_{a}/2c}-\frac{\theta_{5-k}(2{\bf{q}}_{a}+\eta)}{\theta_{1}(2{\bf{q}}_{a})}\,e^{-{\bf{p}}_{a}/2c}\Big)\,,\quad k=1,2,3\,.}\end{array}$ (25)
The Poisson brackets between ${\bf S}_{k}^{a}(\eta)$ computed through $\{{\bf{p}}_{a},{\bf{q}}_{b}\}=\delta_{ab}$ reproduce the classical Sklyanin algebra [14]:
$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i}(\eta),{\bf S}^{a}_{j}(\eta)\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}(\eta){\bf S}^{a}_{k}(\eta)\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0}(\eta),{\bf S}^{a}_{i}(\eta)\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}(\eta){\bf S}^{a}_{k}(\eta)\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,,}\end{array}$ (26)
@@
where the half-periods $\omega_{j}$ (of the underlying elliptic curve with moduli $\tau$)
@@
are numerated according to numeration of the Pauli matrices. The expressions ${\bf S}^{a}_{k}(\eta)$ (25) are the classical analogues of the representation of quantum Sklyanin algebra through difference operators (the shift operators become exponents of momenta in classical mechanics). The limit $\eta\to 0$ is regular for the functions ${\bf S}^{a}_{k}(\eta)$ (25), and the structure constants of the Sklyanin algebra (26) are independent of $\eta$. Therefore, expressions ${\bf S}^{a}_{k}$ (23) have the same Poisson brackets:
$\begin{array}[]{c}\displaystyle{c\{{\bf S}^{a}_{i},{\bf S}^{a}_{j}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{0}{\bf S}^{a}_{k}\,,}\\ \\ \displaystyle{c\{{\bf S}^{a}_{0},{\bf S}^{a}_{i}\}=-\imath\varepsilon_{ijk}{\bf S}^{a}_{j}{\bf S}^{a}_{k}\big(\wp(\omega_{j})-\wp(\omega_{k})\big)\,.}\end{array}$ (28)
See details in [13] and [20].
The Sklyanin algebra (26) is generated by the classical quadratic exchange relation for the Lax matrices (19)
$\begin{array}[]{c}\displaystyle{\{{\mathcal{L}}_{1}^{a}(z,\eta),{\mathcal{L}}_{2}^{b}(w,\eta)\}=\frac{\delta^{ab}}{c}\,[{\mathcal{L}}_{1}^{a}(z,\eta){\mathcal{L}}_{2}^{a}(w,\eta),r_{12}(z-w)]}\end{array}$ (29)
with the elliptic $r$-matrix
$\begin{array}[]{c}\displaystyle{r_{12}(z-w)=\frac{1}{2}\sum\limits_{k=1}^{3}\varphi_{k}(z-w)\sigma_{k}\otimes\sigma_{k}\,.}\end{array}$ (30)
The center of the algebra (29) (and therefore, center of (25) or (26)) is generated by
$\begin{array}[]{c}\displaystyle{\det{{\mathcal{L}}}^{a}(z,\eta)={\bf C}_{2}^{a}(\eta)-\wp(z){\bf C}_{1}^{a}(\eta)}\end{array}$ (31)
with the Casimir functions
$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}(\eta)=\big({\bf S}^{a}_{1}(\eta)\big)^{2}+\big({\bf S}^{a}_{2}(\eta)\big)^{2}+\big({\bf S}^{a}_{3}(\eta)\big)^{2}\,,}\\ \\ \displaystyle{{\bf C}_{2}^{a}(\eta)=\big({\bf S}^{a}_{0}(\eta)\big)^{2}+\sum\limits_{k=1}^{3}\big({\bf S}^{a}_{k}(\eta)\big)^{2}\wp(\omega_{k})\,.}\end{array}$ (32)
The representation (25) yields
$\begin{array}[]{c}\displaystyle{{\bf C}_{1}^{a}(\eta)=\Big(\frac{\theta_{1}(\eta)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,,\qquad{\bf C}_{2}^{a}(\eta)=\wp(\eta)\Big(\frac{\theta_{1}(\eta)}{\theta_{1}^{\prime}(0)}\Big)^{2}\,.}\end{array}$ (33)
The representation (23) appears from (25) in the limit $\eta\to 0$. Then for ${\bf C}_{1}^{a}=({\bf S}^{a}_{1})^{2}+({\bf S}^{a}_{2})^{2}+({\bf S}^{a}_{3})^{2}$ and ${\bf C}_{2}^{a}=({\bf S}^{a}_{0})^{2}+\sum\limits_{k=1}^{3}({\bf S}^{a}_{k})^{2}\wp(\omega_{k})$ defined through the functions (23) we have
@@
Then from (31) we also conclude that for the Lax matrices ${{\mathcal{L}}}^{a}(z)$ (19)
$\begin{array}[]{c}\displaystyle{\det{{\mathcal{L}}}^{a}(z)=\det{{\mathcal{L}}}^{a}(z,\eta)\big|_{\eta=0}=1\,.}\end{array}$ (35)